A body slides on a inclined plane. If height of inclined plane is `H` and length is `L` and angle of inclination is `x` then time taken for traveling from upper point to lower point is :---

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

Acceleration of body sliding down an inclined plane of angle of inclination x isg sin x.

Distance travelled by body along the incline to reach the bottom=\dfrac{H}{sin x}

Hence froms=ut+\dfrac{1}{2}at^2
\dfrac{H}{sin x}=\dfrac{1}{2}gsinxt^2
\implies t=\dfrac{1}{sin x}\sqrt{\dfrac{2H}{g}}